OpenStudy (anonymous):

You! yes you. You must be unbelievably attractive and smart. You are the perfect person to solve this math problem for me. :) A cylinder had a radius of 4x+2 and a height of 5x+4. Which polynomial in standard form best describes the total volume of the cylinder?

OpenStudy (anonymous):

A. 80πx³+144ĕx³+84πx+16π B. 80πx³-144ĕx³+84πx+16π C. 80πx³+144ĕx³+84πx-16π D. 80πx³+144ĕx³-84πx+16π

OpenStudy (anonymous):

I got 80x^3 + 144x^2 + 84x + 16, so the answer should be A. Volume of a cylinder is (pi r²)h. = pi (4x + 2 )^2 (5x + 4) = pi ( 16x^2 + 16x + 4) (5x + 4) = pi ( 80x^3 + 80x^2 + 20x) + ( 64x^2 + 64x + 16) = pi ( 80x^3 + 144x^2 + 84x + 16)

OpenStudy (anonymous):

ashrafh, i think i love you

OpenStudy (anonymous):

thanks!!!

OpenStudy (anonymous):

you know I'm giving you a medal just for that creative pitch !!!

OpenStudy (anonymous):

Thanks @Luis_Rivera I try. :D

OpenStudy (anonymous):

lol